OpenStudy (anonymous):

Find the solution set of 2x^4-9x+11x^2-4x-6=0, if one of the roots is 1-i.

OpenStudy (anonymous):

the other one is 1+1

OpenStudy (anonymous):

so this thing factors as \[ (x-(1+i))(x-(1-i))Q(x)\]

OpenStudy (anonymous):

first two terms multiply to give \[x^2-2x+2\]

OpenStudy (anonymous):

so your annoying job is to divide \[2x^4-9x^3+11x^2-4x-6\] by \[x^2-2x+2\]

OpenStudy (anonymous):

first off is it clear that \[(x-(1+i))(x-(i-i))=x^2-2x+2\]?

OpenStudy (anonymous):

yessss :)

OpenStudy (anonymous):

good. because it is easy once you know how to do it. -1-1 = -2 and 1 + 1 = 2

OpenStudy (anonymous):

now you have a choice

OpenStudy (anonymous):

you can use long division, or you can think

OpenStudy (anonymous):

if you want to use long division i cannot show that here. but i can help you think through it if you like

OpenStudy (anonymous):

would it work with synthetic?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

because you are not dividing by x - r

OpenStudy (anonymous):

you are dividing by a quadratic.

OpenStudy (anonymous):

so thinking is really the best method

OpenStudy (anonymous):

ohhhh i see /:

OpenStudy (anonymous):

i will start you out. \[2x^4 -9x^3+11x^2-4x-6=(x^2-2x+2)(ax^2+bx+c)\]

OpenStudy (anonymous):

now a and c are obvious yes?

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

name them

OpenStudy (anonymous):

A is 2 and c is 11?

OpenStudy (anonymous):

a is right, c is wrong. think about what the constant must be

OpenStudy (anonymous):

is it -6?

OpenStudy (anonymous):

nope but getting warmer. \[c\times 2=-6\]

OpenStudy (anonymous):

that is the only way you get the constant from this product

OpenStudy (anonymous):

so -3?

OpenStudy (anonymous):

clear yes because all the other terms will have x's in them

OpenStudy (anonymous):

so c must be -2 to give -6 as the constant and of course a = 2 so you get \[2x^4\]

OpenStudy (anonymous):

so you just factored it, right?

OpenStudy (anonymous):

yes. now comes the only hard part, finding b \[2x^4-9x^3+11x^2-4x-6=(x^2-2x+2)(2x^2+bx-3)\]

OpenStudy (anonymous):

but we have many choices to find it, and it doesn't matter which method we pick

OpenStudy (anonymous):

for example to get the "x" term it will come from \[2\times bx -3\times -2x\] \[=(2b+6)x\] and we know that we have -4x so \[2b+6=-4\]

OpenStudy (anonymous):

i am just looking at the product and seeing what will give me an x term. we know we have to end up with -4x and if you multiply on the right you will have it

OpenStudy (anonymous):

is it -5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can also check with the \[x^3\] term

OpenStudy (anonymous):

we know we have to end up with \[-9x^3\] and that term will come from \[bx^3-4x^3\]

OpenStudy (anonymous):

so we know that \[b-4=-9\] again telling us b = -5

OpenStudy (anonymous):

oohh makes sense!!

OpenStudy (anonymous):

it is more annoying to do it with the x^2 terms because there are several ways to get x^2

OpenStudy (anonymous):

and of course you only have to check with one because (assuming your arithmetic is correct) it HAS to work because we know it factors.

OpenStudy (anonymous):

i think this is easier than long division, but you can do it that way if you like. it is a mess. again the idea is to think what the first and last term has to be, and then find the middle one

OpenStudy (anonymous):

like that myininaya?

OpenStudy (anonymous):

thhats so much batter than long division. i just need to get this down. but thank you so much!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

you can do polynomial division with more than just a linear term